OpenStudy (anonymous):

The endpoints of a movable rod (3 ft length) have coordinates (x,0) and (0,y) the position of the end on the x-axis is x(t)=2sin(πt/6) where t is the time in seconds , find the time of one complete cycle of the rod.

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@hartnn

OpenStudy (sidsiddhartha):

 so u have the equation x^2+y^2=1 u have x(t)=2sin(πt/6) completing one cycle means the rod have to get back from where it's started so it need to have the same x and y coordinates so time needed=(2*pi)/(pi/6)=12 sec